# The ubiquity of variable radio emission and spin-down rates in pulsars

### v2

@@ Abstract
Pulsars are often lauded for their (relative) rotational and radio emission stability over long time scales. However, long-term observing programmes are identifying an increasing number of pulsars that deviate from this preconceived notion. Using Gaussian process regression and Bayesian inference techniques, we investigated the emission and rotational stability of 259 radio pulsars that have been monitored using Murriyang, the Parkes 64 m radio telescope, over the past three decades. We found 238 pulsars that display significant variability in their spin-down rates, 52 of which also exhibit changes in profile shape. Including 23 known state-switching pulsars, this represents the largest catalogue of variable pulsars identified to date and indicates that these behaviours are ubiquitous among the wider population. The intensity of spin-down fluctuations positively scales with increasing pulsar spin-down rate, with only a marginal dependence on spin-frequency. This may have substantial implications for ongoing searches for gravitational waves in the ensemble timing of millisecond pulsars. We also discuss challenges in explaining the physical origins of quasi-periodic and transient profile/spin-down variations detected among a subset of our pulsars.
@@
Mode or state switching between two or more discrete radio emission states was identified soon after the discovery of the first pulsars, manifesting as changes in the observed intensity, or shape or position in pulse longitude of the pulse profile components (see Backer 1970b, c; Lyne 1971). It is a broadband phenomenon, having been detected both at radio and X-ray wavelengths (Hermsen et al., 2013, 2017, 2018), and takes place over a variety of timescales. This can be as short as a one or several rotations of a neutron star, or slow secular variations that take place over many months or years. In some cases the emission appears to briefly switch off and back on again (nulls; Backer 1970a) or remain off for extended periods of time (intermittency; Kramer et al. 2006). Long-term monitoring experiments found that the timing of some pulsars displayed quasi-periodic processes with a range of apparent periods (e.g. Cordes & Greenstein 1981; Cordes & Downs 1985; Stairs et al. 2000). Both of the fluctuations in radio emission and spin-down rate were eventually linked, initially through observations of the intermittent pulsar PSR B193124 (Kramer et al., 2006), and later through increasingly large samples of pulsars that switch between two or more quasi-stable spin-down levels or undergo transient spin-down events (Kramer et al., 2006; Hobbs et al., 2010; Lyne et al., 2010; Brook et al., 2014, 2016; Shaw et al., 2022). Other studies have revealed pulsars with variations in either spin-down or pulse shape (and even polarization) that appear uncorrelated with one another (Brook et al., 2016; Shaw et al., 2022; Basu et al., 2024), though this may be a selection effect imparted by limited telescope sensitivity and observing cadence.
@@
Exactly what mechanism drives these behaviours, and whether they originate from the interior dynamics or magnetospheric fluctuations of neutron stars, is presently unknown. Free precession of a non-axisymmetric neutron star was once a popular hypothesis for explaining highly periodic changes in pulsar profile shapes and spin-down rates (Stairs et al., 2000). However, the discovery that such spin-down rate changes were linked to short-term emission variations has cast doubt on whether free precession is the true, or only, mechanism behind these behaviours (Lyne et al., 2010; Jones, 2012; Kerr et al., 2016; Stairs et al., 2019). Simultaneous emission switching at radio and X-ray wavelengths point toward a magnetospheric reconfiguration being responsible for at least some of the mode-switching behaviour (Hermsen et al., 2013). Emission and spin-down variations that seemingly coincided with glitches in several pulsars (e.g, Weltevrede et al. 2011; Keith et al. 2013) may point to crustal events such as starquakes that couple the internal dynamics of the neutron star to the magnetosphere (Akbal et al., 2015; Yuan et al., 2021). Indeed the 2016 glitch in the Vela pulsar (PSR J08354510/B083345) was associated with a short-lived quenching of its radio emission (Palfreyman et al., 2018), speculated to be due to a quake-induced magnetic disturbance (Bransgrove et al., 2020). External triggers such as interactions between pulsars and minor bodies are another popular hypothesis. These include interactions with individual in-falling asteroids (Brook et al., 2014), debris disks (Li, 2006; Shannon et al., 2013; Jennings et al., 2020b), and impacts from interstellar objects (Pham et al., 2024). In such cases, the ionised remains of a wayward asteroid alter the plasma content of the pulsar magnetosphere, subsequently enhancing or attenuating the radio emission and altering the spin-down state. Many previous attempts to explain pulsar variability focussed on either individual objects or small samples with similar characteristics (e.g., Brook et al., 2014, 2016; Lower et al., 2023; Basu et al., 2024). Understanding whether one or more of these processes are responsible for these profile/spin-down variations requires population-level studies of a large ensemble of pulsars.
@@
Determining the prevalence of pulsar rotation/emission state changes across the population is vital to building a complete picture of the processes that contribute to the overall noise budget of a pulsar. Knowing how these behaviours scale as functions of spin frequency and spin-down rate is particularly important for pulsar timing array experiments that utilise an ensemble of millisecond pulsars to search for nanohertz-frequency gravitational waves. Even the most stable millisecond pulsars have been found to display profile shape variations and emission state switching that can hamper such experiments (Shannon et al., 2016; Brook et al., 2018; Goncharov et al., 2021; Miles et al., 2022; Nathan et al., 2023; Jennings et al., 2024). It may also bring us a step closer to determining the origins of unexpected time-dependent changes in the common noise process seen by pulsar timing array experiments (see Reardon et al. 2023).
@@
In this work, we present the largest sample of pulsars found to display emission and spin-down variations to date, by applying the Gaussian process regression techniques of (Brook et al., 2016; Brook et al., 2018) to observations collected over the past three decades by Murriyang, the CSIRO Parkes 64 m radio telescope. In Section 2 we briefly describe the data collection process. Section 3 details the methodologies used for obtaining updated pulsar timing solutions and modelling the pulse profile shape and spin-down variations, the results of which are presented in Section 4. We discuss how these effects vary across the population and the implications for pulsar timing array experiments in Section 5. Finally, we summarise the results and make concluding statements in Section 6.
2 Observations
@@
The pulsar data analysed in this work were collected using Murriyang between 2007-2023 under the ‘young’ pulsar timing (P574) programme (Weltevrede et al., 2010; Johnston et al., 2021). Our sample covers 259 isolated, non-recycled, rotation-powered pulsars with characteristic ages ranging from 1.6 kyr (PSR J15135908) to 0.4 Gyr (PSR J08204114). In Figure 1 we show the positions of our pulsar sample in period-period-derivative space alongside those of the broader ATNF Pulsar Catalogue (Manchester et al., 2005)111https://www.atnf.csiro.au/research/pulsar/psrcat/. Note, there is a slight bias towards pulsars with high rotational kinetic energy loss rates due to the programme’s continued support of the Fermi mission (Smith et al., 2023). Pulsar observations were undertaken with the 10/50222Later re-tuned to cover the 40-cm band due to increased interference., 20-cm Multibeam, H-OH and Ultra-Wideband Low (UWL) receivers (Granet et al., 2001; Staveley-Smith et al., 1996; Granet et al., 2011; Hobbs et al., 2020) and saved to psrfits format archives (Hotan et al., 2004; van Straten et al., 2012). Details of the specific observing set-up, radio-frequency excision, and calibration steps can be found in Namkham et al. (2019); Parthasarathy et al. (2019); Lower et al. (2021); Johnston et al. (2021). Where possible, we also made use of extended timing datasets collected using legacy signal processors going back to the 1990s, which are available in the Parkes Observatory Pulsar Data Archive (Hobbs et al., 2011)333https://data.csiro.au/domain/atnf. Pulse times of arrival (ToAs) were obtained following the prescription outlined previously in Parthasarathy et al. (2019), Lower et al. (2021) and Lower et al. (2023). We averaged the individual observations in time and frequency to form one-dimensional pulse profiles, which were then cross-correlated against a noise-free template using the Fourier-domain Monte Carlo method built into the pat tool in psrchive. Note that we employed observing-band specific templates for each pulsar.
@@ 3.1 Updated pulsar timing
A substantial fraction of the pulsars required updates to their timing models over those published in the ATNF Pulsar Catalogue. Phase coherent solutions were derived following the manual procedures outlined in Parthasarathy et al. (2019) and Lower et al. (2021), where pulsar positions, proper motions, spin frequency (), spin-down () and on occasion, second spin-frequency derivative (), were re-fit using a combination of tempo2 and temponest (Hobbs et al., 2006; Lentati et al., 2014). The majority of pulsars that required inclusion of were those previously studied in Parthasarathy et al. (2020) and Lower et al. (2021). Glitch events were identified by eye as discontinuities or sharp features in the timing residuals. These were corrected for by fitting for the corresponding changes in and using tempo2. Obtaining accurate measurements of the recovery parameters is challenging and imperfectly removed recovery signals will introduce unwanted artefacts in our spin-down analysis. A detailed analysis of the new glitch events and overall updates to the timing of these pulsars is beyond the scope of this work, and will be reported elsewhere. Once a coherent solution was obtained, we then applied pulse numbering to the ToAs to maintain an accurate record of every single rotation over our timing baseline. In Figure 2 we show the corresponding timing residuals for our final sample of 259 pulsars.
@@ 3.2 Profile variability maps
In order to model the observed variations in the pulsar profiles and spin-down rates, we followed the general methodology devised by Brook et al. (2016). To detect and then study changes in the radio profile shape, we created profile variability maps that display variations in profile intensity as functions of rotation phase and observation date. Only the total intensity radio profiles collected by the Parkes Multibeam, H-OH and 1216–1600 MHz UWL subband were used in the shape variation analysis. Observations that were visually affected by residual RFI, instrumental artefacts, or displayed an anomalously low signal-to-noise (S/N) ratio were manually flagged and excluded. As in Lower et al. (2023), we performed an initial coarse alignment of the profiles by maximising the correlation between each observation with the highest S/N profile. The alignment was then refined using the iterative scaling and shifting approach outlined in Brook et al. (2018). We created standard templates for each pulsar by computing the median normalised flux within each phase bin, which were then subtracted from our observations to produce profile residuals. These profile residuals were then fitted with non-parametric models on a per phase bin basis via Gaussian process regression to produce profile variability maps. Variability maps where the absolute peak phase-bin standard deviation () was greater than three times the median off-pulse root-mean-square often indicated the presence of substantial profile shape variability above random noise fluctuations in the data. However, this metric is sensitive to distortions in the data caused by residual RFI, hence visual inspection of the variability maps and profiles was used to filter out these outliers. We also note that apparent shape changes due to the different central frequencies of the multibeam (2007-2016, 2017-2019), H-OH (2016-2017) and UWL (2019 onwards) receivers limits our sensitivity to more subtle temporal shape changes.
@@ 3.3 Modelling spin-down rate variations
where and refer to the minimum and maximum inferred spin-down derived from the Gaussian process regression, and is the mean spin-down uncertainty computed via equations 9 and 10 of Brook et al. (2016). We used a threshold of for defining when a pulsar displayed substantial spin-down variability. To test the level of correlation between any observed variations in profile shape and spin-down rate, we computed the Spearman rank correlation coefficient between each phase bin of the profile variability map and for time-lags ranging between half the length of the per-pulsar observing span (Brook et al., 2016; Shaw et al., 2022; Lower et al., 2023).
@@ 4 Results and analysis
Our variability analysis identified 238 pulsars that displayed significant () changes in over time. We present the resulting spin-down timeseries for the 238 variable pulsars in Figure 5, and list the recovered Gaussian process hyper-parameters and inferred rotational properties in Tables 2 and 3 in Appendix A. In general, we find the largest changes in spin-down can be broadly described by four different categories:
@@
Profile variability maps were generated for 214 of pulsars that displayed significant variations, which can be found in the Supplementary Materials. An example of one is shown in Figure 8 for PSR J10485832, where the upper and lower panels present the varying spin spin-down rate and profile shape changes respectively. The remaining 24 of the 238 variable pulsars had insufficient per-epoch to generate accurate variability maps. From these maps, we found that 52 pulsars showed substantial changes in profile shape. We also identified several pulsars that appear to show intense epoch-to-epoch shape variations across the entire profile that are approximately symmetric about the peak. This behaviour can be attributed to profile jitter, resulting from the finite number of rotations that are recorded in our relatively short, per-pulsar observations. Pulsars that were most strongly affected by this include PSRs J08374135, J1430-6623 and J1644-4559, which are not included in the profile shape change analysis.
@@ 4.1 Known variable pulsars
This 0.33 s pulsar that has been reported to switch between two emission states: ‘mode A’ where the emission occurs slightly earlier in pulse phase, and ‘mode B’ where it is both weaker and arrives slightly later (Zhang et al., 2020). Our spin-down model displays rapid, short-duration decreases in spin-down rate with a quasi-periodic spacing of 268 d. These spin-down events typically last between 130-200 d with typical amplitudes of %. The profile variability map shows the trailing edge of the profile displays changes on an similarly rapid timescale, usually appearing as an excess of emission whenever the pulsar is in the lower spin-down state that is only mildly correlated with the spin-down timeseries. An excess of emission in this part of the profile would point to the pulsar preferentially emitting in ‘mode B’ between the quasi-periodic decreases in spin-down rate.

### v3

@@ Abstract
Pulsars are often lauded for their (relative) rotational and radio emission stability over long time scales. However, long-term observing programmes are identifying an increasing number of pulsars that deviate from this preconceived notion. Using Gaussian process regression and Bayesian inference techniques, we investigated the emission and rotational stability of 259 isolated radio pulsars that have been monitored using Murriyang, the Parkes 64 m radio telescope, over the past three decades. We found that 238 pulsars display significant variability in their spin-down rates, 52 of which also exhibit changes in profile shape. Including 23 known state-switching pulsars, this represents the largest catalogue of variable pulsars identified to date and indicates that these behaviours are ubiquitous among the wider population. The intensity of spin-down fluctuations positively scales with increasing pulsar spin-down rate, with only a marginal dependence on spin-frequency. This may have substantial implications for ongoing searches for gravitational waves in the ensemble timing of millisecond pulsars. We also discuss challenges in explaining the physical origins of quasi-periodic and transient profile/spin-down variations detected among a subset of our pulsars.
@@
Mode or state switching between two or more discrete radio emission states was identified soon after the discovery of the first pulsars, manifesting as changes in the observed intensity, or shape or position in pulse longitude of the pulse profile components (see Backer 1970b, c; Lyne 1971). It is a broadband phenomenon, having been detected both at radio and X-ray wavelengths (Hermsen et al., 2013, 2017, 2018), and takes place over a wide range of timescales. This can be as short as one or several rotations of a neutron star, or slow secular variations that take place over many months or years. In some cases the emission appears to briefly switch off and back on again (nulls; Backer 1970a) or remain off for extended periods of time (intermittency; Kramer et al. 2006). Long-term monitoring experiments found that the timing of some pulsars displayed quasi-periodic processes with a range of apparent periods (e.g. Cordes & Greenstein 1981; Cordes & Downs 1985; Stairs et al. 2000). Fluctuations in both radio emission and spin-down rate were eventually linked, initially through observations of the intermittent pulsar PSR B193124 (Kramer et al., 2006), and later through increasingly large samples of pulsars that switch between two or more quasi-stable spin-down levels or undergo transient spin-down events (Kramer et al., 2006; Hobbs et al., 2010; Lyne et al., 2010; Brook et al., 2014, 2016; Shaw et al., 2022). Other studies have revealed pulsars with variations in either spin-down or pulse shape (and even polarization) that appear uncorrelated with one another (Brook et al., 2016; Shaw et al., 2022; Basu et al., 2024), though this may be a selection effect imparted by limited telescope sensitivity and observing cadence.
@@
Exactly what mechanism drives these behaviours, and whether they originate from the interior dynamics or magnetospheric fluctuations of neutron stars, is presently unknown. Free precession of a non-axisymmetric neutron star was once a popular hypothesis for explaining highly periodic changes in pulsar profile shapes and spin-down rates (Stairs et al., 2000). However, the discovery that such spin-down rate changes were linked to short-term emission variations has cast doubt on whether free precession is the true, or only, mechanism behind these behaviours (Lyne et al., 2010; Jones, 2012; Kerr et al., 2016; Stairs et al., 2019). Simultaneous emission switching at radio and X-ray wavelengths point toward a magnetospheric reconfiguration being responsible for at least some of the mode-switching behaviour (Hermsen et al., 2013). Emission and spin-down variations that seemingly coincided with glitches in several pulsars (e.g, Weltevrede et al. 2011; Keith et al. 2013) may point to crustal events such as starquakes that couple the internal dynamics of the neutron star to the magnetosphere (Akbal et al., 2015; Yuan et al., 2021). Indeed, the 2016 glitch in the Vela pulsar (PSR J08354510/B083345) was associated with a short-lived quenching of its radio emission (Palfreyman et al., 2018), speculated to be due to a quake-induced magnetic disturbance (Bransgrove et al., 2020). External triggers such as interactions between pulsars and minor bodies are another popular hypothesis. These include interactions with individual in-falling asteroids (Brook et al., 2014), debris disks (Li, 2006; Shannon et al., 2013; Jennings et al., 2020b), and impacts from interstellar objects (Pham et al., 2024). In such cases, the ionised remains of a wayward minor body alter the plasma content of the pulsar magnetosphere, subsequently enhancing or attenuating the radio emission and altering the spin-down state. Many previous attempts to explain pulsar variability focussed on either individual objects or small samples with similar characteristics (e.g., Brook et al., 2014, 2016; Lower et al., 2023; Basu et al., 2024). Understanding whether one or more of these processes are responsible for these profile/spin-down variations requires population-level studies of a large ensemble of pulsars.
@@
Determining the prevalence of pulsar rotation/emission state changes across the population is vital to building a complete picture of the processes that contribute to the overall noise budget of a pulsar. Knowing how these behaviours scale as functions of spin frequency and spin-down rate is particularly important for pulsar timing array experiments that utilise an ensemble of millisecond pulsars to search for nanohertz-frequency gravitational waves (Manchester et al., 2013; Manchester & IPTA, 2013; Kramer & Champion, 2013; McLaughlin, 2013). Even the most stable millisecond pulsars have been found to display profile shape variations and emission state switching that can hamper such experiments (Shannon et al., 2016; Brook et al., 2018; Goncharov et al., 2021; Miles et al., 2022; Nathan et al., 2023; Jennings et al., 2024). It may also bring us a step closer to determining the origins of unexpected time-dependent changes in the common noise process seen by pulsar timing array experiments (see Reardon et al. 2023).
@@
In this work, we present the largest sample of pulsars found to display emission and spin-down variations to date, by applying the Gaussian process regression techniques of (Brook et al., 2016; Brook et al., 2018) to observations collected over the past three decades by Murriyang, the CSIRO Parkes 64 m radio telescope. In Section 2 we briefly describe the data collection and reduction process. Section 3 details the methodologies used for obtaining updated pulsar timing solutions and modelling the pulse profile shape and spin-down variations, the results of which are presented in Section 4. We discuss how these effects vary across the population and the implications for pulsar timing array experiments in Section 5. Finally, we summarise the results and make concluding statements in Section 6.
2 Observations and data reduction
@@
The pulsar data analysed in this work were collected using Murriyang between 2007-2023 under the ‘young’ pulsar timing (P574) programme (Weltevrede et al., 2010; Johnston et al., 2021). Our sample covers 259 isolated, non-recycled, rotation-powered pulsars with characteristic ages ranging from 1.6 kyr (PSR J15135908) to 0.4 Gyr (PSR J08204114). In Figure 1 we show the positions of our pulsar sample in period-period-derivative space alongside those of the broader ATNF Pulsar Catalogue (Manchester et al., 2005)111https://www.atnf.csiro.au/research/pulsar/psrcat/. Note that there is a slight bias towards pulsars with high rotational kinetic energy loss rates due to the programme’s continued support of the Fermi mission (Smith et al., 2023). Pulsar observations were undertaken with the 10/50222Later re-tuned to cover the 40-cm band due to increased interference., 20-cm Multibeam, H-OH and Ultra-Wideband Low (UWL) receivers (Granet et al., 2001; Staveley-Smith et al., 1996; Granet et al., 2011; Hobbs et al., 2020) and saved to psrfits format archives (Hotan et al., 2004; van Straten et al., 2012). Details of the specific observing set-up, radio-frequency excision, and calibration steps can be found in Namkham et al. (2019); Parthasarathy et al. (2019); Lower et al. (2021); Johnston et al. (2021). Where possible, we also made use of extended timing datasets collected using legacy signal processors going back to the 1990s, which are available in the Parkes Observatory Pulsar Data Archive (Hobbs et al., 2011)333https://data.csiro.au/domain/atnf.
Pulse times of arrival (ToAs) were obtained following the prescription outlined previously in Parthasarathy et al. (2019), Lower et al. (2021) and Lower et al. (2023). We averaged the individual observations in time and frequency to form one-dimensional pulse profiles, which were then cross-correlated against a noise-free template using the Fourier-domain Monte Carlo (FDM) method built into the pat tool in psrchive (Hotan et al., 2004; van Straten et al., 2012). Note that we employed observing band-specific templates for each pulsar.
@@ 3.1 Updated pulsar timing
A substantial fraction of the pulsars required updates to their timing models over those published in the ATNF Pulsar Catalogue. Phase coherent solutions were derived following the manual procedures outlined in Parthasarathy et al. (2019) and Lower et al. (2021), where pulsar positions, proper motions, spin frequency (), spin-down rate () and on occasion, second spin-frequency derivative (), were re-fit using a combination of tempo2 and temponest (Hobbs et al., 2006; Lentati et al., 2014). The majority of pulsars that required inclusion of were those previously studied in Parthasarathy et al. (2020) and Lower et al. (2021). Glitch events were identified by eye as discontinuities or sharp features in the timing residuals. These were corrected for by fitting the corresponding changes in and using tempo2. Obtaining accurate measurements of the recovery parameters is challenging and imperfectly removed recovery signals will introduce unwanted artefacts in our spin-down analysis. A detailed analysis of the new glitch events and overall updates to the timing of these pulsars is beyond the scope of this work, and will be reported elsewhere. Once a coherent solution was obtained, we then applied pulse numbering to the ToAs to maintain an accurate record of every single rotation over our timing baseline. In Figure 2 we show the corresponding timing residuals for our final sample of 259 pulsars.
@@ 3.2 Profile variability maps
In order to model the observed variations in the pulsar profiles and spin-down rates, we followed the general methodology devised by Brook et al. (2016). To detect and then study changes in the radio profile shape, we created profile variability maps that display variations in profile intensity as functions of rotational phase and observation date. Only the total intensity radio profiles collected by the Parkes Multibeam, H-OH and 1216–1600 MHz UWL subband were used in the shape variation analysis. Observations that were visually affected by residual RFI, instrumental artefacts, or displayed an anomalously low signal-to-noise (S/N) ratio were manually flagged and excluded. As in Lower et al. (2023), we performed an initial coarse alignment of the profiles by maximising the correlation between each observation with the highest S/N profile. The alignment was then refined using the iterative scaling and shifting approach outlined in Brook et al. (2018). We created standard templates for each pulsar by computing the median normalised flux within each phase bin, which were then subtracted from our observations to produce profile residuals. These profile residuals were then fitted with non-parametric models on a per phase bin basis via Gaussian process regression to produce profile variability maps. Variability maps where the absolute peak phase-bin standard deviation () was greater than three times the median off-pulse root-mean-square often indicated the presence of substantial profile shape variability above random noise fluctuations in the data. However, this metric is sensitive to distortions in the data caused by residual RFI, hence visual inspection of the variability maps and profiles was used to filter out these outliers. We also note that apparent shape changes due to the different central frequencies of the multibeam (2007-2016, 2017-2019), H-OH (2016-2017) and UWL (2019 onwards) receivers limits our sensitivity to more subtle temporal shape changes.
@@ 3.3 Modelling spin-down rate variations
where and refer to the minimum and maximum inferred absolute spin-down rates derived from the Gaussian process regression, and is the mean spin-down uncertainty computed via equations 9 and 10 of Brook et al. (2016). We used a threshold of for defining when a pulsar displayed substantial spin-down variability. To test the level of correlation between any observed variations in profile shape and spin-down rate, we computed the Spearman rank correlation coefficient between each phase bin of the profile variability map and for time-lags ranging between half the length of the per-pulsar observing span (Brook et al., 2016; Shaw et al., 2022; Lower et al., 2023).
@@ 4 Results and analysis
Our variability analysis identified 238 pulsars that displayed significant () changes in over time. We present the resulting spin-down timeseries for the 238 variable pulsars in Figure 5, and list the recovered Gaussian process hyper-parameters and inferred rotational properties in Tables 2 and 6 in Appendix A. In general, we find the largest changes in spin-down can be broadly described by four different categories:
@@
Profile variability maps were generated for 214 pulsars that displayed significant variations, which can be found in the Supplementary Materials. One such example is shown in Figure 8 for PSR J10485832, where the upper and lower panels present the varying spin spin-down rate and profile shape changes respectively. The remaining 24 of the 238 variable pulsars had insufficient per-epoch to generate accurate variability maps. From these maps, we found that 52 pulsars showed substantial changes in profile shape. We also identified several pulsars that appear to show intense epoch-to-epoch shape variations across the entire profile that are approximately symmetric about the peak. This behaviour can be attributed to profile jitter, resulting from the finite number of rotations that are recorded in our relatively short, per-pulsar observations. Pulsars that were most strongly affected by this include PSRs J08374135, J1430-6623 and J1644-4559, which are not included in the profile shape change analysis.
@@ 4.1 Known variable pulsars
This 0.33 s pulsar that has been reported to switch between two emission states: ‘mode A’ where the emission occurs slightly earlier in pulse phase, and ‘mode B’ where it is both weaker and arrives slightly later (Zhang et al., 2020). Our spin-down model displays rapid, short-duration decreases in spin-down rate with a quasi-periodic spacing of 268 d. These spin-down events typically last between 130-200 d with typical amplitudes of %. The profile variability map shows the trailing edge of the profile displays changes on a similarly rapid timescale, usually appearing as an excess of emission whenever the pulsar is in the lower spin-down state that is only mildly correlated with the spin-down timeseries. An excess of emission in this part of the profile would point to the pulsar preferentially emitting in ‘mode B’ between the quasi-periodic decreases in spin-down rate.